# A Fractional Calculus Framework for Open Quantum Dynamics: From Liouville to Lindblad to Memory Kernels

### v1

@@ IV.1 Why Markovian dephasing misses $e^{-Q(t)}$?
Since the entire Markovian kernel is a $\delta$-function concentrated at the present time, this is different from the exact kernel (29) that possesses a long tail inherited from the bath spectrum $J(\omega)$. Therefore, the time-homogeneous Lindbladian cannot capture the long-time memory that produces $u(t)=e^{-Q(t)}$. The “missing terms” are precisely the nonlocal (finite-memory) contributions. It is worth noting that, in order to reproduce $u(t)=e^{-Q(t)}$, one can choose the rate $\gamma$ to be time-dependent. For example, with $\gamma(t)=\tfrac{1}{2}\dot{Q}(t)$, we have
@@
This more flexible time-local dephasing model merely re‑encodes the exact memory in $\gamma(t)$ through $Q(t)$ to reproduce the exact dephasing law. Moreover, if $\gamma(t)$ were to become temporarily negative (possible for structured environments), the dynamics is non-Markovian (not CP-divisible) Rivas2014; Hall2014, and the GKSL form with constant positive rates is no longer valid.
@@
Figure 3: Comparison of exact coherence from Figure 2 with Markovian and non-Markovian fits. The asymptotic propagation of the coherence amplitude $|\langle\sigma_{+}(t)\rangle|$ for three different bath exponents $\chi=0.5$, $1.0$, and $1.5$ are computed by constant $\gamma$ (Markovian, left panel) and time-local $\gamma(t)=\tfrac{1}{2}\dot{Q}(t)$ (non-Markovian, right panel), respectively. The constant rate $\gamma$ is obtained by fitting $\log|\langle\sigma_{+}(t)\rangle|$ to a linear decay in the intermediate-time regime $t\in[2/\omega_{c},60/\omega_{c}]$ (blue shade, left panel), while solid lines are obtained by numerically integrating Eq. (26).
@@
The analysis above highlights that a constant-rate Lindbladian cannot capture the long-tailed memory encoded in $Q(t)$, while a time-dependent rate merely re-parametrizes this memory without providing structural insight. A more fundamental generalization is therefore to embed the memory directly into the generator through fractional differentiation. We next show that replacing the first-order derivative in the Lindblad equation with a fractional derivative produces a compact, CPTP, and physically transparent model that reproduces the exact non-Markovian coherence $e^{-Q(t)}$ across spectral regimes.
Figure 4: Asymptotic propagation of the coherence amplitude $|\langle\sigma_{+}(t)\rangle|$ for the spin–boson model obtained from the single–order fractional propagation and compared with the numerically exact result (solid lines, computed by direct integration of Eq. (26)). Results are shown for three spectral exponents, $\chi=0.5$, $1.0$, and $1.5$. The fractional parameters $(\alpha,\lambda)$ are determined by the least–squares fitting to the exact propagation within the indicated intermediate–time windows (blue shades, left panels): $t\!\in\![2/\omega_{c},20/\omega_{c}]$ (top), $t\!\in\![2/\omega_{c},40/\omega_{c}]$ (middle), and $t\!\in\![2/\omega_{c},60/\omega_{c}]$ (bottom). Circle markers denote the fractional results; the shaded regions mark the fitting intervals.
IV.2 Single–order fractional Lindbladian for structured baths
@@
where $(\alpha,\lambda)$ is a scalar pair that can be optimized for reproducing the expected dephasing. The same Mittag–Leffler envelope has recently appeared in the context of the generalized fractional Rabi problem, where fractional differentiation modulates the coherent oscillations of a driven two-level system lopez2025generalisedfractionalrabiproblem.
@@
Figure 4 compares the coherence amplitude $|\langle\sigma_{+}(t)\rangle|$ obtained from the single–order fractional propagation (40) with the numerically exact NZ evolution for representative bath exponents $\chi=0.5$, $1.0$, and $1.5$. For each case, the fractional parameters $(\alpha,\lambda)$ are optimized within the shaded intermediate–time windows and then the parameters are fixed for the fractional propagation over the entire time domain. Concretely, $(\alpha,\lambda)$ are obtained by minimizing the discrete root–mean–square deviation between the fractional coherence $u_{\alpha}(t;\alpha,\lambda)$ and the NZ result $u_{\mathrm{NZ}}(t)$ over a fitting interval $t_{\mathrm{start}},t_{\mathrm{end}}]$, cf. Appendix D. The fitting windows are chosen to be several times the bath correlation time $\tau_{B}\sim 1/\omega_{c}$ (so that both the initial curvature and the onset of the long–time tail are resolved) yet shorter than the recurrence time set by the high–frequency cutoff.
@@
This consistent fitting procedure is applied to the sub–Ohmic, Ohmic, and super–Ohmic examples shown in the top, middle, and bottom panels, respectively. As can be seen, the fractional model accurately reproduces the exact coherence for the sub–Ohmic bath (with the deviation <0.02<0.02 over the entire time domain regardless of the selected time windows for the parameter fitting), demonstrating that a single fractional order can faithfully mimic the memory kernel encoded in $Q(t)$. Since $Q(t)$ is completely determined by the bath correlation function $C(t)=\langle B(t)B(0)\rangle$ and hence by the spectral density $J(\omega)$, the optimized $(\alpha,\lambda)$ can be viewed as compact, two–parameter summaries of the influence of $C(t)$ over the chosen time window. For the Ohmic and super–Ohmic baths, the agreement gradually worsens as $\chi$ increases, reflecting the reduced influence of long–tailed correlations. On the other hand, the accuracy in the Ohmic baths can be systematically improved by fitting over a broader time window, which allows the optimized $(\alpha,\lambda)$ to capture both the short–time curvature and the slower asymptotic decay of the exact solution. Nevertheless, improvements for the super–Ohmic bath are modest compared to the other two baths, and its large residual deviation (>0.05>0.05) motivates the modified fractional ansatz, which incorporates the finite plateau $u_{\infty}$ characteristic of super-Ohmic environments.
@@
$u(t)=u_{\infty}+(1-u_{\infty})\,E_{\alpha}(-\lambda t^{\alpha})\quad\Rightarrow\quad v(t)=\frac{u(t)-u_{\infty}}{1-u_{\infty}}=E_{\alpha}(-\lambda t^{\alpha}).$ (41)
This affine form bounds the coherence between $u_{\infty}$ and 1 and ensures the correct long–time plateau. We therefore fit the plateau–normalized curve $v(t)$ within the shaded intermediate–time window to obtain $(\alpha,\lambda)$, and then reconstruct $u(t)$ from Eq. (41). The optimized fractional orders are systematically larger than s1s-1, compensating for the overly rapid decay of the fixed–order model and capturing the slower relaxation characteristic of structured baths. As shown in Figure 5, this plateau–anchored fractional ansatz accurately reproduces both the intermediate–time curvature and the long–time coherence plateau for all super–Ohmic exponents considered ($\chi=1.2$, $1.5$, and $1.8$).
@@
Although $(\alpha,\lambda)$ can be determined by least–squares fitting, a simple physics–based choice can avoid nonlinear optimization entirely. From the sub–Ohmic kernel, the Laplace-domain relation $\mathcal{L}\{t^{-\beta}\}\!\sim\!s^{\beta-1}$ implies that a convolution kernel $t^{-\beta}$ entering the memory term $\int_{0}^{t}\!(t-\tau)^{-\beta}u(\tau)\,d\tau$ is equivalent to a fractional derivative of order $1-\beta$ Podlubny1999; Kilbas2006. Hence, the asymptotic behavior $Q^{\prime}(t)\!\sim\!t^{-\chi}$ yields the scaling rule $\alpha\simeq 1-\chi$, which ensures that a fractional convolution reproduces the long–memory tail of the exact $Q(t)$. In practice, when the accessible simulation window lies in the pre–asymptotic regime, the best effective order is slightly larger ($\alpha_{\mathrm{eff}}\!\in(1-\chi,1)$), capturing the stretched–exponential envelope of $u_{\mathrm{NZ}}(t)$ before the Mittag–Leffler power law sets in. A quick, data–driven estimate can be obtained from the local log–log slope of $X(t)=-\ln|u_{\mathrm{NZ}}(t)|$ (for sub-Ohmic and Ohmic baths) or $X(t)=-\ln|v(t)|$ (for super-Ohmic bath) as $\alpha_{\text{loc}}(t_{i})=\big[\ln X(t_{i+1})-\ln X(t_{i-1})\big]/\big[\ln t_{i+1}-\ln t_{i-1}\big],$ with $\alpha$ chosen as the median over the region where the slope is approximately constant. Once $\alpha$ is fixed, $\lambda$ follows from a single point in the same window by solving Eq. (40) or (41). This optimization–free selection rule recovers the expected scaling $\alpha\!\approx\!1-\chi$ for sub–Ohmic baths and provides reliable initial guesses for the least–squares refinement used in Figs. 4 and 5.
@@
IV.3 Unified Interpretation and Quantum Algorithms for Fractional Dynamics
The fractional framework is able to provide both a unified interpretation of non-Markovian quantum dynamics and a constructive pathway toward their quantum simulation. Numerical benchmarks on the spin–boson model show that a stationary single-order fractional Lindbladian can closely track the exact Nakajima–Zwanzig coherence over broad time windows, with errors at the few-percent level. Operationally, the fractional order $\alpha$ continuously tunes the degree of memory: lowering $\alpha$ slows relaxation, broadens correlation times, and induces algebraic envelopes, all consistent with the long-tail structure inherited from the bath spectral density $J(\omega)$. Within this hierarchy, the Liouville, Lindblad, and fractional master equations correspond respectively to unitary, Markovian, and structured non-Markovian evolution. Fractional differentiation introduces memory directly at the generator level, embedding power-law temporal correlations without violating complete positivity, and providing a continuous interpolation between reversible ($\alpha\!=\!1$) and strongly non-Markovian ($\alpha\!\ll\!1$) limits.

### v2

@@ III.1 Why Markovian dephasing misses $e^{-Q(t)}$?
Since the entire Markovian kernel is a $\delta$-function concentrated at the present time, this is different from the exact kernel (33) that possesses a long tail inherited from the bath spectrum $J(\omega)$. Therefore, the time-homogeneous Lindbladian cannot capture the long-time memory that produces $u(t)=e^{-Q(t)}$. The left panel of Figure 5 explicitly illustrates the discrepancy, where a constant Markovian rate $\gamma$ produces a purely exponential decay that systematically deviates from the exact coherence $u(t)=e^{-Q(t)}$. The “missing terms” are precisely the nonlocal (finite-memory) contributions. It is worth noting that, in order to reproduce $u(t)=e^{-Q(t)}$, one can choose the rate $\gamma$ to be time-dependent. For example, with $\gamma(t)=\tfrac{1}{2}\dot{Q}(t)$, we have
@@
Numerically, as can be seen from the right panel of Figure 5, a time-local rate $\gamma(t)=\tfrac{1}{2}\dot{Q}(t)$ can reproduce the exact behavior only by encoding the full non-Markovian memory directly into the time-dependent rate. This more flexible time-local dephasing model merely re‑encodes the exact memory in $\gamma(t)$ through $Q(t)$ to reproduce the exact dephasing law. Moreover, if $\gamma(t)$ were to become temporarily negative (possible for structured environments), the dynamics is non-Markovian (not CP-divisible) Rivas2014; Hall2014, and the GKSL form with constant positive rates is no longer valid.
@@
Figure 5: Comparison of exact coherence from Figure 4 with Markovian and non-Markovian fits. The asymptotic propagation of the coherence amplitude $|\langle\sigma_{+}(t)\rangle|$ for three different bath exponents $\chi=0.5$, $1.0$, and $1.5$ are computed by constant $\gamma$ (Markovian, left panel) and time-local $\gamma(t)=\tfrac{1}{2}\dot{Q}(t)$ (non-Markovian, right panel), respectively. The constant rate $\gamma$ is obtained by fitting $\log|\langle\sigma_{+}(t)\rangle|$ to a linear decay in the intermediate-time regime $t\in[2/\omega_{c},60/\omega_{c}]$ (blue shade, left panel), while solid lines are obtained by numerically integrating Eq. (30). The parameters shown here ($\gamma$ and $\gamma(t)$) belong to Markovian and time-local models only and are not related to the fractional parameters $(\alpha,\lambda)$ introduced in Figure 7.
@@
We emphasize that the quantities fitted in Figure 5 (the constant rate $\gamma$ or the time-local rate $\gamma(t)$) pertain solely to the Markovian and time-local Lindblad descriptions. The associated analysis highlights that a constant-rate Lindbladian cannot capture the long-tailed memory encoded in $Q(t)$, while a time-dependent rate merely re-parametrizes this memory without providing structural insight. The purpose of Figure 5 is therefore to demonstrate the limitations of these Lindbladian descriptions and to motivate the introduction of a more fundamental generalization that embeds the memory directly into the generator. We next show that replacing the first-order derivative in the Lindblad equation with a fractional derivative produces a compact, CPTP, and physically transparent model that reproduces the exact non-Markovian coherence $e^{-Q(t)}$ across spectral regimes. Remarkably, the parameters used in the Lindbladian descriptions in this section are not related to the fractional model introduced in the next section and do not involve the fractional order $\alpha$.
Figure 6: Comparison between the exact coherence $u(t)=e^{-Q(t)}$ (solid lines) and the parameter–free fractional predictions $u_{\alpha}(t)=E_{\alpha}(-\lambda t^{\alpha})$ (circles) for sub–Ohmic ($\chi=0.5$), Ohmic ($\chi=1$), and super–Ohmic ($\chi=1.5$) baths. The fractional parameters $(\alpha_{\rm pred},\lambda_{\rm pred})$ used here are obtained directly from microscopic information: (Left) $\alpha=2$ and $\lambda=2A_{\chi}$ follow from the universal short-time Gaussian behavior of $Q(t)$ ($t<0.2/\omega_{c}$); (Right) the long-time values of $(\alpha_{\rm pred},\lambda_{\rm pred})$ follow from the asymptotic structure of $Q(t)$ determined by the spectral density $J(\omega)\propto\omega^{\chi}$ ($5/\omega_{c}\leq t\leq 10^{5}/\omega_{c}$). The analytic fractional model reproduces both the correct short-time curvature and the correct long-time decay or plateau behavior for all bath types without any fitting. As expected, the discrepancies appear in the intermediate-time crossover regime, where neither the Gaussian expansion nor the asymptotic form of $Q(t)$ is accurate; this region reflects subleading features of the true spectral density and may be used to refine or infer additional structure in $J(\omega)$.
III.2 Single–order fractional Lindbladian for structured baths
@@
where $(\alpha,\lambda)$ do not correspond to the Markovian or time-local rates $(\gamma,\gamma(t))$ discussed previously, and can be found for reproducing the expected dephasing. The same Mittag–Leffler envelope has recently appeared in the context of the generalized fractional Rabi problem, where fractional differentiation modulates the coherent oscillations of a driven two-level system lopez2025generalisedfractionalrabiproblem.
The rest of this section is organized as follows: (i) a physics-based recipe for predicting $(\alpha,\lambda)$ from the spectral density; (ii) an optional refinement scheme (least-squares fitting) used only for benchmarking against the exactly solvable spin–boson model; and (iii) a dedicated super–Ohmic ansatz required when the coherence saturates at a finite plateau.
Table 2: Predictive fractional parameters $(\alpha_{\rm pred},\lambda_{\rm pred})$ obtained directly from microscopic inputs. The universal short-time parameters follow from the Gaussian expansion of $Q(t)$, while the long-time parameters follow from the asymptotic form of $Q(t)$ determined by the bath exponent $\chi$. All derivations are given in Appendix D.
Regime / Bath type Structure of $Q(t)$ $\alpha_{\rm pred}$     $\lambda_{\rm pred}$
[…]
Sub–Ohmic ($0<\chi<1$) $C_{\chi}\,t^{1-\chi}$ $1-\chi$     $C_{\chi}\Gamma(2-\chi)$
Ohmic ($\chi=1$) $\tfrac{2\eta}{\pi}\,\ln t$ $\eta$     $1/\Gamma(1-\eta)$
Super–Ohmic ($\chi>1$) $Q_{\infty}-D_{\chi}\,t^{-(\chi-1)}$ $\chi-1$     $\frac{e^{Q_{\infty}}-1}{D_{\chi}\Gamma(2-\chi)}$
III.2.1 Predicting $(\alpha,\lambda)$ directly from $J(\omega)$
A major advantage of the fractional formulation is that $(\alpha,\lambda)$ can be specified directly from microscopic inputs, which are the bath exponent $\chi$ and the asymptotic structure of the dephasing kernel $Q(t)$ (or the structure of spectral density $J(\omega)$). Appendix D provides the full derivations for all three bath classes (sub–Ohmic, Ohmic, super–Ohmic). The resulting predictive rules are summarized in Table 2 and are used verbatim in Figure 6. As can be seen from the figure, in the short-time regime ($t<0.2/\omega_{c}$ the dephasing kernel exhibits the universal Gaussian expansion $Q(t)\approx A_{\chi}t^{2}$, which fixes the fractional short-time parameters to $\alpha_{\rm pred}=2$ and $\lambda_{\rm pred}=2A_{\chi}$. While our main fractional master equation is formulated for $0<\alpha<1$, this short-time assignment $\alpha_{\rm pred}=2$ should be understood as an effective Caputo order in the scalar coherence equation used solely to match the initial Gaussian curvature; it is not employed as a full fractional Lindblad evolution for all times. Beyond this window, the coherence is governed entirely by the long-time asymptotics of $Q(t)$, and the fractional model transitions to the memory exponent $\alpha_{\rm pred}$ listed in the table. These exponents reproduce the correct non-Markovian decay or saturation behavior over the full long-time interval ($5/\omega_{c}\leq t\leq 10^{5}/\omega_{c}$).
The accuracy of these analytic predictions is also illustrated in Figure 6, which compares the parameter-free fractional solutions with exact results for all three bath types. The agreement at early and late times is a direct consequence of the fact that both regimes are controlled by analytic structures of $Q(t)$, and ultimately of $J(\omega)$: the short-time curvature is inherited from the Gaussian prefactor $A_{\chi}$, while the long-time behavior reflects the power-law (or plateau-anchored) asymptotics of $Q(t)$ determined by the bath exponent $\chi$. As expected, the fractional prediction deviates from the exact dynamics in the intermediate-time interval ($5/\omega_{c}\lesssim t\lesssim 10^{4}/\omega_{c}$). This is precisely the regime where neither the short-time Gaussian form nor the long-time asymptotic expansion of $Q(t)$ is accurate, and where subleading features of the true spectral density $J(\omega)$ play a dominant role. The fact that the discrepancy is confined to this intermediate region suggests that the fractional model may serve as a sensitive diagnostic of missing structure in $J(\omega)$: improving the bath model in this time window would sharpen agreement across all times. Indeed, one may either refine the assumed form of $J(\omega)$ or, conversely, use partial fitting in the intermediate-time range to infer additional information about the underlying spectral density. We explore this perspective in the following subsection.
Overall, the comparison demonstrates that the fractional model can be parameterized predictively and entirely from microscopic inputs, with the intermediate-time deviations providing useful guidance on how to refine or learn the spectral structure of the environment.
Figure 7: Asymptotic propagation of the coherence amplitude $|\langle\sigma_{+}(t)\rangle|$ for the spin–boson model obtained from the single–order fractional propagation and compared with the numerically exact result (solid lines, computed by direct integration of Eq. (30)). Results are shown for three spectral exponents, $\chi=0.5$, $1.0$, and $1.5$. The fractional parameters $(\alpha,\lambda)$ are determined by the least–squares fitting to the exact propagation within the indicated intermediate–time windows (blue shades, left panels): $t\!\in\![2/\omega_{c},20/\omega_{c}]$ (top), $t\!\in\![2/\omega_{c},40/\omega_{c}]$ (middle), and $t\!\in\![2/\omega_{c},60/\omega_{c}]$ (bottom). Circle markers denote the fractional results; the shaded regions mark the fitting intervals. The parameters $(\alpha,\lambda)$ shown here are fractional parameters and should not be compared with the Markovian or time-local rates in Figure 5.
III.2.2 Refinement via least–squares fitting (benchmarking and bath inference)
To quantify the accuracy of the single-order fractional model and to extract an effective memory exponent in the intermediate regime where discrepancies emerge, we refine $(\alpha_{\rm pred},\lambda_{\rm pred})$ by least-squares fitting to the exact spin–boson coherence. This refinement is used strictly as a diagnostic and benchmarking tool: it reveals how the fractional parameters summarize bath-induced memory over a specific temporal interval and provides a principled route to inferring additional structure in $J(\omega)$. Accordingly, we focus on the interval $5/\omega_{c}\leq t\leq 10^{2}/\omega_{c}$, where the deviation between the predictive fractional model and the exact solution in Figure 7 is most pronounced. Extending the comparison into the later-time regime ($10^{2}/\omega_{c}$$10^{4}/\omega_{c}$) does not change the qualitative conclusions, as the deviation decreases monotonically once the dynamics enters the asymptotic regime captured by the analytic structure of $Q(t)$.
@@ III.2.2 Refinement via least–squares fitting (benchmarking and bath inference)
Figure 7 compares the coherence amplitude $|\langle\sigma_{+}(t)\rangle|$ obtained from the single–order fractional propagation (44) with the numerically exact NZ evolution for representative bath exponents $\chi=0.5$, $1.0$, and $1.5$. For each case, The fractional parameters $(\alpha,\lambda)$ are optimized within the shaded intermediate–time windows and then the parameters are fixed for the fractional propagation over the entire time domain. Concretely, $(\alpha,\lambda)$ are obtained by minimizing the discrete root–mean–square deviation between the fractional coherence $u_{\alpha}(t;\alpha,\lambda)$ and the NZ result $u_{\mathrm{NZ}}(t)$ over a fitting interval $t_{\mathrm{start}},t_{\mathrm{end}}]$, cf. Appendix E. The fitting windows are chosen to be several times the bath correlation time $\tau_{B}\sim 1/\omega_{c}$ (so that both the initial curvature and the onset of the long–time tail are resolved) yet shorter than the recurrence time set by the high–frequency cutoff.
@@
This consistent fitting procedure is applied to the sub–Ohmic, Ohmic, and super–Ohmic examples shown in the top, middle, and bottom panels, respectively. As can be seen, the fractional model accurately reproduces the exact coherence for the sub–Ohmic bath (with the deviation <0.02<0.02 over the entire time domain regardless of the selected time windows for the parameter fitting), demonstrating that a single fractional order can faithfully mimic the memory kernel encoded in $Q(t)$. Since $Q(t)$ is completely determined by the bath correlation function $C(t)=\langle B(t)B(0)\rangle$ and hence by the spectral density $J(\omega)$, the optimized $(\alpha,\lambda)$ can be viewed as compact, two–parameter summaries of the influence of $C(t)$ over the chosen time window. For the Ohmic and super–Ohmic baths, the agreement gradually worsens as $\chi$ increases, reflecting the reduced influence of long–tailed correlations. On the other hand, the accuracy in the Ohmic baths can be systematically improved by fitting over a broader time window, which allows the optimized $(\alpha,\lambda)$ to capture both the short–time curvature and the slower asymptotic decay of the exact solution. Nevertheless, improvements for the super–Ohmic bath are modest compared to the other two baths, and its large residual deviation (>0.05>0.05) motivates the modified fractional ansatz as described in the following section, which incorporates the finite plateau $u_{\infty}$ characteristic of super-Ohmic environments.
More importantly, because the exact coherence spans multiple dynamical regimes—Gaussian at short times, stretched-exponential at intermediate times, and algebraic or plateau-anchored at long times—no single exponent can describe the full evolution. Thus a fitted $\alpha$ should be viewed as an effective memory exponent appropriate to the chosen time window, analogous to window-dependent anomalous diffusion exponents in CTRW theory klages2008anomalous; METZLER20001; Metzler_2004; Metzler2014. The variation of fitted $\alpha$ across different windows therefore serves as a quantitative probe of unresolved features in the spectral density $J(\omega)$ and offers a potential route for bath inference. For example, in the sub–Ohmic case, if fitting over an intermediate window yields an effective exponent $\alpha_{\rm fit}$, then matching the stretched-exponential form implies an effective spectral exponent $\chi_{\mathrm{eff}}=1-\alpha_{\rm fit}$. Thus $\alpha_{\rm fit}>\alpha_{\rm pred}=1-\chi$ indicates additional low-frequency weight in $J(\omega)$, e.g., a softened cutoff or a secondary power-law component. We present this only as an illustrative example; developing a systematic reconstruction scheme is left for future work.
III.2.3 Super–Ohmic baths: fractional ansatz with plateau normalization
@@
$u(t)$ $=u_{\infty}+(1-u_{\infty})\,E_{\alpha}(-\lambda t^{\alpha})$
$\Rightarrow~~v(t)$ $=\frac{u(t)-u_{\infty}}{1-u_{\infty}}=E_{\alpha}(-\lambda t^{\alpha}).$ (45)
This affine form bounds the coherence between $u_{\infty}$ and 1 and renders the initial Gaussian decay and most of the intermediate stretched-exponential behavior negligible, thereby isolating the long-time algebraic regime of the dynamics. We therefore fit the plateau–normalized curve $v(t)$ within the shaded intermediate–time window to obtain $(\alpha,\lambda)$, and then reconstruct $u(t)$ from Eq. (45). The optimized fractional orders are systematically larger than s1s-1, compensating for the overly rapid decay of the fixed–order model and capturing the slower relaxation characteristic of structured baths. As shown in Figure 8, this plateau–anchored fractional ansatz accurately reproduces both the intermediate–time curvature and the long–time coherence plateau for all super–Ohmic exponents considered ($\chi=1.2$, $1.5$, and $1.8$). Because the asymptotic tail is mapped to a pure power law, the fractional parameters extracted from $v(t)$ in Figure 8 emphasize the long-time memory exponent and may differ from parameters obtained by fitting the full coherence $u(t)$. For example, for $\chi=1.5$, the fractional order obtained from the plateau-normalized fit ($\alpha=0.62$) is approximately twice that obtained from the full-coherence fit in Figure 7 ($\alpha=0.31$), even when using the same fitting window.
@@
In practice, when the accessible simulation window lies in the pre–asymptotic regime, the best effective order is slightly larger ($\alpha_{\mathrm{eff}}\!\in(1-\chi,1)$), capturing the stretched–exponential envelope of $u_{\mathrm{NZ}}(t)$ before the Mittag–Leffler power law sets in. A quick, data–driven estimate can be obtained from the local log–log slope of $X(t)=-\ln|u_{\mathrm{NZ}}(t)|$ (for sub-Ohmic and Ohmic baths) or $X(t)=-\ln|v(t)|$ (for super-Ohmic bath) as $\alpha_{\text{loc}}(t_{i})=\big[\ln X(t_{i+1})-\ln X(t_{i-1})\big]/\big[\ln t_{i+1}-\ln t_{i-1}\big],$ with $\alpha$ chosen as the median over the region where the slope is approximately constant. Once $\alpha$ is fixed, $\lambda$ follows from a single point in the same window by solving Eq. (44) or (45). This optimization–free selection rule recovers the expected scaling $\alpha\!\approx\!1-\chi$ for sub–Ohmic baths and provides reliable initial guesses for the least–squares refinement used in Figures 7 and 8.
@@
III.3 Unified Interpretation and Quantum Algorithms for Fractional Dynamics
The fractional framework is able to provide both a unified interpretation of non-Markovian quantum dynamics and a constructive pathway toward their quantum simulation. Numerical benchmarks on the spin–boson model show that a stationary single-order fractional Lindbladian can closely track the exact NZ coherence over broad time windows, with errors at the few-percent level. Operationally, the fractional order $\alpha$ continuously tunes the degree of memory: lowering $\alpha$ slows relaxation, broadens correlation times, and induces algebraic envelopes, all consistent with the long-tail structure inherited from the bath spectral density $J(\omega)$. Within this hierarchy, the Liouville, Lindblad, and fractional master equations correspond respectively to unitary, Markovian, and structured non-Markovian evolution. Fractional differentiation introduces memory directly at the generator level, embedding power-law temporal correlations without violating complete positivity, and providing a continuous interpolation between reversible ($\alpha\!=\!1$) and strongly non-Markovian ($\alpha\!\ll\!1$) limits.